# “Where am I?” Scene Retrieval with Language

### v1

@@ 4.1 Datasets
For each scene in our ScanScribe dataset, there are on average 20 text descriptions, leading to a total of 4472 descriptions that describe 218 unique scenes. Descriptions in this ScanScribe dataset are generated using GPT-3 [16], a text template, and an input of the scene graph annotations. The text descriptions also do not always describe the scene to its entirety, as the authors first sample varying sizes of subgraphs from the scene graphs, and then generate a description. Examples of these text descriptions and their corresponding text graphs can be found in supp. m. We separate this ScanScribe dataset into a training set and a testing set, where the testing set has 1116 descriptions from 55 unique scenes. This split will be made public for full reproducibility.
Human-Annotated Dataset. We ask human annotators to observe sequences of images from a scene within the 3RScan dataset [74], and then write a short description of what they see, focusing on objects and spatial relationships. These scenes already have corresponding scene graphs from the 3DSSG [75] dataset. The human annotators are asked to describe the scene in a natural, “human-like” way. We collect 147 natural language descriptions of 142 unique scenes in this way. We consider this human-annotated dataset as another testing set and refer to it from now on as the “Human dataset.” Examples of these text descriptions and their corresponding text graphs can be found in Figure 2.
Table 2: The recall values for matching the top $k$ scenes, out of all scenes, to our text-queries. All scene candidates are from the ScanScribe (top) [83] and the collected Human (bottom) datasets. Together with the baselines, Text2Pos and CLIP2CLIP, we show three versions of our proposed method, “match-prob” and “cos-sim” employing different scoring functions, and “ret-based” obtaining matches efficiently by pre-computing the scene graph embeddings in an offline stage and running simple nearest-neighbors search. Note that our three versions are all trained with the matching probability and cosine similarity in the objective function. They differ in how they calculate the matching score.
@@ 4.2 Evaluation Setup
For all mentioned models, we first train with a validation set, separate from the training and testing set, and then use the validation loss to determine a suitable number of epochs. Then we retrain the models on the entire training set.
Text2Pos. We compare our method to the “coarse-localization” module from Text2Pos [38]. This module performs coarse-localization in a large-scale outdoor point cloud using a natural language query. Text2Pos does this by separating a large scale point cloud into “cells” and identifies the top $k$ most likely “cells.” The “cells” and text-query are first individually embedded by a point cloud embedding model and a text embedding model respectively, and then matched according to their cosine similarity score. Since Text2Pos takes point cloud inputs, we extract the point clouds from the corresponding scenes in our datasets and consider each scene as a single “cell.” Notably, other concurrent work, Text2Loc [79] and RET [76] build off of Text2Pos, but their models are not available. We compare against a fine-tuned and retrained version of the “coarse-localization” module from Text2Pos, trained on a training set from ScanScribe dataset.
@@
CLIP2CLIP. A second baseline we compare to is one that we implement ourselves, and we call it CLIP2CLIP. The idea is as follows: we embed the text-queries and images from the scenes using CLIP [58], and then attempt to match text description to scenes via these CLIP embeddings. The context size of CLIP’s text encoder only supports lengths up to 77, which is why we first separate the sentences into phrases delimited by commas or periods. Then we encode each smaller phrase, and take the average over all the phrases as the “text-query encoding.” For the images, we simply encode each image using CLIP’s image encoder. Then, in order to get a matching score between a “text-query” and a scene of images, we calculate the cosine similarity between the text encoding and all image encodings from the scene, and take the maximum score as the matching score between that text and scene. During inference, given a “text-query,” we encode it and calculate the cosine similarity scores between the text encoding and image encodings of each scene. The highest scoring pair is the match.
Text2SGM. We show three versions of the joint embedding model from our Text2SGM pipeline. The first (match-prob) uses the “matching probability” outputs from the model as the matching scores between a text-query and a scene graph. The second one (cos-sim) calculates a cosine similarity score between the text-query and a scene graph embeddings produced by the model.
The third version differs from the first two in one key aspect: it first individually embeds the scene graph into embedding vectors. Here we separately embed the text-query and the scene graph, and we calculate the cosine similarity between their embeddings. Given that our model requires an input pair of a text-query and a scene graph, we investigate how different text-queries affect the resulting scene graph embeddings and vice versa. Detailed results of this investigation are in the supplementary material, but we observe that the resulting embeddings do not have a high variance while varying the other paired input. As such, for the third version of our model, we fix the opposing paired input with the same text-query or scene graph. This text-query or scene graph is randomly sampled from the training dataset, and we embed all inputs with this fixed counterpart. In essence, this is retrieval-based (red-based) matching between the text-query and scene graph embeddings.
It is worth noting that all three versions are trained with both matching probability and cosine similarity as a part of the objective function. We differentiate these three models by how the “similarity score” is calculated, with either the matching probability, or the cosine similarity.
@@ 4.3 Metrics
We measure two accuracy metrics of the different methods. The first measures a “top k out of 10” accuracy. We try to match one randomly sampled text-query against 10 scene graphs, out of which 1 scene graph is the correct one. We measure if the correct pair is matched within the top 1, 2, 3, and 5 matches out of 10. Similarly, we measure a “top k out of the entire dataset” accuracy, where we try to retrieve the correct graph from the entire dataset. We report accuracy values in Table 1 and Table 2. In addition to accuracy, we also measure the inference time of matching 1 text-query to a scene graph. We report these times in 3. Lastly, for each method, we calculate the size of storing our datasets in memory. More specifically, we calculate the embedding size of each method, since these scene embeddings can be precomputed and stored for later retrieval.
@@ 4.4 Results
Here, we first present the recall values for the three versions of our model compared against baselines on the ScanScribe and the captured Human datasets [83].
Table 1 reports the recalls when performing top-$k$ selection from $10$ candidates scenes. It also shows that our method outperforms Text2Pos as well as CLIP2CLIP by a large margin on both datasets in almost all configurations. Only in the Human dataset, for the top 5 out of 10 and the top 75 out of the entire dataset, does CLIP2CLIP outperform our method. For the most part, using the embedding cosine similarity as the scoring mechanism performs the best. We posit that this is due to the high-dimensional joint embedding capturing the nuances in the graphs, something not achieved by using just the matching probability (i.e., a single real number). We additionally recognize that the requirement for a “known-words” mapping in Text2Pos severely limits its ability to generalize to the open-set text-queries from the Human dataset. Furthermore, these results show that Text2Pos does not generalize to indoor scenes.
Our retrieval-based pipeline performs almost as well as, and in some instances better than, the pipeline that jointly calculates the embeddings. This suggests our joint embedding model learns an adequately meaningful embedding. Also, we observe that the CLIP2CLIP performance is much more competitive to our performance on the Human dataset, compared to the ScanScribe dataset. Seeing as our training data is from the same distribution as the ScanScribe dataset, our much better performance here is to be expected. However, given that CLIP [58] is pre-trained on a much larger-scale dataset, we also see that it generalizes better on the open-set text-queries from the Human dataset. Nevertheless, these recall values suggest that our method successfully retrieves scenes given a text-query describing the scene.
Results when performing top-$k$ selection from all candidates scenes are reported in Table 2. We can observe very similar trends as when selecting from $10$ scenes. Text2Pos is on the level of random selection for both scenes in all recall values. The proposed Text2SGM method improves upon them significantly, achieving meaningful results across almost all datasets and configurations. The highest accuracy is achieved, again, by employing the cosine similarity. The retrieval-based approach does not lag much behind in terms of accuracy, suggesting that the obtained graph embeddings are insensitive to the input text and, thus, generalize well.
@@ 4.6 Ablation Studies
Table 4 reports the recall values together with their standard deviations on both tested datasets, employing cosine similarity ($\theta$), matching score ($m$) or both ($\theta,m$), in the loss objective, when training the proposed network. Interestingly, while training with only the matching probability lags behind other losses on the ScanScribe dataset, it leads to the best performance on the Human dataset. Conversely, a cosine similarity loss alone leads to the most accurate results on ScanScribe, while being marginally less accurate on the Human dataset. Employing both losses leads to the best results by a large margin.
Table 5 reports the results when varying the number of self and cross-attention modules in the proposed method. The best results are achieved with only one module, with every additional one significantly reducing the accuracy. This could be due to over-smoothing problem [21, 63] as previously mentioned. Thus, we use a single module in all experiments.
Table 6 demonstrates the effect of using different numbers of attention heads on the two tested datasets. While there is no clear trend when increasing the number, minor improvements can still be observed across the tested configurations. These improvements manifest in marginally increasing recall and decreasing standard deviation.
@@
Table 6: Number of Attention Heads Ablation. We show recall values for adding different numbers of attention heads during training. These are recall values for the ScanScribe and Human datasets, at a top $k$ out of 10 precision.
Top k ScanScribe Human
out of 10 1 2 3 5 1 2 3 5
1 $62.04\pm 2.31$ $79.65\pm 1.93$ $87.27\pm 1.46$ $94.28\pm 0.99$ $50.53\pm 2.42$ $71.70\pm 2.30$ $82.44\pm 1.84$ $92.01\pm 1.28$
2 $63.09\pm 2.37$ $81.39\pm 1.78$ $88.68\pm 1.43$ $95.63\pm 0.90$ $50.46\pm 2.49$ $71.70\pm 2.03$ $82.61\pm 1.73$ $92.79\pm 1.37$
4 $62.02\pm 2.13$ $79.47\pm 1.73$ $87.17\pm 1.54$ $94.85\pm 1.05$ $50.35\pm 2.68$ $72.00\pm 2.33$ $82.16\pm 2.13$ $91.34\pm 1.41$
6 $63.89\pm 2.14$ $81.56\pm 1.75$ $88.79\pm 1.48$ $95.34\pm 1.02$ $50.75\pm 2.38$ $72.65\pm 2.11$ $83.79\pm 1.91$ $93.17\pm 1.23$

### v2

@@ 4.1 Datasets
For each scene in the ScanScribe dataset, there are on average 20 text descriptions, leading to a total of 4472 descriptions that describe 218 unique scenes. Descriptions in this ScanScribe dataset are generated using GPT-3 [14]. The text descriptions also do not always describe the scene to its entirety, as the authors first sample varying sizes of subgraphs from the scene graphs, and then generate a description. Examples of these text descriptions and their corresponding text graphs can be found in supp. m. We separate this ScanScribe dataset into a training set and a testing set, where the testing set has 1116 descriptions from 55 unique scenes. This split will be made public for full reproducibility.
Human-Annotated Dataset. We ask human annotators to observe sequences of images from a scene within the 3RScan dataset [66], and then write a short description of what they see, focusing on objects and spatial relationships. These scenes already have corresponding scene graphs from the 3DSSG [67] dataset. The human annotators are asked to describe the scene in a natural, “human-like” way. We collect 147 natural language descriptions of 142 unique scenes in this way. We consider this human-annotated dataset as another testing set and refer to it from now on as the “Human dataset.” These descriptions were not generated from the scene-graph, thus there is no bias. Examples of these text descriptions and their corresponding text graphs can be found in Figure 2.
Table 2: The recall values for matching the top $k$ scenes, out of all scenes, to our text-queries. All scene candidates are from the ScanScribe (top) [74] and the collected Human (bottom) datasets. Together with the baselines, Text2Pos and CLIP2CLIP, we show three versions of our pipeline, “match-prob” and “cos-sim” employing different matching score calculations, and “ret-based” (precomputing scene graph embeddings and doing retrieval-based matching). Note that the underlying joint embedding models in all three models are trained with the matching probability and cosine similarity in the loss function. They differ in how they calculate the matching score.
@@ 4.2 Evaluation Setup
For all mentioned models, we first train on a training set. We also have a validation set that we use to determine the best model parameters. We then retrain our model on the combined training and validation set to maximize the amount of training data. Finally, we test on completely separate testing sets.
Text2Pos. We compare our method to the “coarse-localization” module from Text2Pos [39]. This module performs coarse-localization in a large-scale outdoor point cloud using a natural language query. Text2Pos does this by separating a large scale point cloud into “cells” and identifies the top $k$ most likely “cells.” The “cells” and text-query are first individually embedded by a point cloud embedding model and a text embedding model respectively, and then matched according to their cosine similarity. Since Text2Pos takes point cloud inputs, we extract the point clouds from the corresponding scenes in our datasets and consider each scene as a single “cell.” Notably, other concurrent work, Text2Loc [70] and RET [68] build off of Text2Pos, but their models are not available. We compare against a fine-tuned and retrained version of the “coarse-localization” module from Text2Pos, trained on a training set from ScanScribe dataset.
@@
CLIP2CLIP. A second baseline we compare to is one that we implement ourselves, and we call it CLIP2CLIP. The idea is as follows: we embed the text-queries and images from the scenes using CLIP [50], and then attempt to match text description to scenes via these CLIP embeddings. The context size of CLIP’s text encoder only supports lengths up to 77, which is why we first separate the sentences into phrases delimited by commas or periods. Then we encode each smaller phrase, and take the average over all the phrases as the “text-query encoding.” For the images, we simply encode each image using CLIP’s image encoder. Then, in order to get a matching score between a “text-query” and a scene of images, we calculate the cosine similarity between the text encoding and all image encodings from the scene, and take the maximum score as the matching score between that text and scene. During inference, given a “text-query,” we encode it and calculate the cosine similarity between the text encoding and image encoding of each scene. The highest scoring pair is the match. The CLIP model is not fine-tuned or retrained.
Text2SGM. We present two procedures for calculating the matching score. The first (match-prob) uses the “matching probability” outputs from the model as the matching scores between a text-query and a scene graph. The second one (cos-sim) calculates a cosine similarity between the text-query and scene graph embedding produced by the model. Both procedures use the same underlying joint embedding model.
We also present a third retrieval-based procedure for finding text and scene graph matches. Given that our pipeline requires an input pair of a text-query and a scene graph–due to the cross attention module–we investigate how different text-queries affect the resulting scene graph embeddings and vice versa. Detailed results of this investigation are in the supplementary material, but we observe that the embeddings are still consistently identifiable, despite being paired with different inputs. As such, for the third version of our model, we fix the opposing paired input using the same text-query or scene graph. This text-query or scene graph is randomly sampled from the training dataset, and we embed all inputs with this fixed counterpart. Given a text-query, we pass it to the model along with a fixed scene graph input, and get the text-graph embedding. Then we find the highest cosine similarity match in a database of pre-computed scene graph embeddings. We call this our retrieval-based (ret-based) matching approach.
To reiterate, the underlying joint embedding model is the same across all three “matching score” calculation procedures. The model is trained with both the matching probability and cosine similarity loss as a part of the objective function. The “match-prob” procedure uses the matching probability as the matching score. The “cos-sim” procedure uses the cosine similarity between embeddings as the matching score. The “ret-based” procedure also uses the cosine similarity as the matching score, but pre-computes a database of scene graph embeddings.
@@ 4.3 Metrics
We measure two recall metrics, the first measures a “top k out of 10” recall. We try to match one randomly sampled text-query against 10 scene graphs, out of which 1 scene graph is the correct one. We measure if the correct pair is matched within the top 1, 2, 3, and 5 matches out of 10. Similarly, we measure a “top k out of the entire dataset” recall, where we try to retrieve the correct graph from the entire dataset. We report recall values in Table 1 and Table 2. In addition to recall, we also measure the inference time of matching 1 text-query to a scene graph. We report these times in Table 3. Lastly, for each method, we calculate the size of storing our datasets in memory. More specifically, we calculate the embedding size of each method, since these scene embeddings can be precomputed (ret-based method) and stored for later retrieval.
@@ 4.4 Results
We first present the recall values for the three versions of our pipeline compared against baselines on the ScanScribe and the captured Human datasets [74].
Table 1 reports the recalls when performing top-$k$ selection from $10$ candidates scenes. It also shows that our method outperforms Text2Pos as well as CLIP2CLIP by a large margin on both datasets in almost all configurations. Only in the Human dataset, for the top 5 out of 10 and the top 75 out of the entire dataset (Table 2), does CLIP2CLIP outperform our method. For the most part, using the embedding cosine similarity as the scoring mechanism performs the best. We posit that this is due to the high-dimensional joint embedding capturing the nuances in the graphs, something not achieved by using just the matching probability (i.e., a single real number). We additionally recognize that the requirement for a “known-words” mapping in Text2Pos severely limits its ability to generalize to the open-set text-queries from the Human dataset. Furthermore, these results show that Text2Pos does not generalize to indoor scenes.
Our retrieval-based pipeline performs almost as well as, and in some instances better than, the pipeline (cos-sim) that jointly calculates the embeddings. This suggests our joint embedding model learns an adequately meaningful embedding. Also, we observe that the CLIP2CLIP performance is much more competitive to our performance on the Human dataset, compared to the ScanScribe dataset. Seeing as our training data is from the same distribution as the ScanScribe dataset, our much better performance here is to be expected. However, given that CLIP [50] is pre-trained on a much larger-scale dataset, we also see that it generalizes better on the open-set text-queries from the Human dataset. Nevertheless, these recall values suggest that our method successfully retrieves scenes given a text-query describing the scene.
Results when performing top-$k$ selection from all candidates scenes are reported in Table 2. Text2Pos performs on the level of random selection for both scenes in all recall values. The proposed Text2SGM method improves upon this significantly, achieving meaningful results across almost all datasets and configurations. The highest recall is achieved, again, by employing the cosine similarity as the scoring mechanism. The retrieval-based approach does not lag much behind in terms of recall, suggesting that the obtained graph embeddings are meaningful despite using the same input text-query, and thus, generalize well. Additionally, we show results in Table 4 using text-queries that are generated by GPT-4 [1] from images of the 3DSSG scenes. Even though we are now biasing towards the CLIP2CLIP method since we generate our text-queries from images, we still perform more accurately.
Table 4: The recall values from using text-queries generated by GPT-4 from images of the 3DSSG scenes. We generate text-queries from images because the ScanScribe test set in Tables 1 and 2 has queries generated from the scene graphs. The Human dataset, in addition to this experiment, eliminates that bias. We measure top-k out of 10 recall on the ScanScribe test set, note that we generate 1 text-query per scene.
Top k 1 2 3 5
CLIP2CLIP $33.07\pm 0.21$ $52.52\pm 0.23$ $66.00\pm 0.21$ $\textbf{82.78}\pm 0.13$
Text2SGM (cos-sim) $\textbf{34.22}\pm 1.77$ $\textbf{56.67}\pm 1.70$ $\textbf{68.78}\pm 2.35$ $82.11\pm 1.23$
@@ 4.6 Ablation Studies
Table 5 reports the recall values together with their standard deviations on both tested datasets, employing cosine similarity ($\theta$), matching probability ($m$) or both ($\theta,m$), in the loss objective, when training the proposed network. Interestingly, while training with only the matching probability lags behind other losses on the ScanScribe dataset, it leads to the best performance on the Human dataset. Conversely, a cosine similarity loss alone leads to the most accurate results on ScanScribe, while being marginally less accurate on the Human dataset. Employing both losses leads to the best results by a large margin.
Table 6 reports the results when varying the number of self and cross-attention modules in the proposed method. The best results are achieved with only one module, with every additional one significantly reducing the recall. This could be due to over-smoothing problem [18, 55] as previously mentioned. Thus, we use a single module in all experiments.
@@ 5 Conclusion
Acknowledgements. This work was partially funded by the ETH RobotX research grant, the Hasler Stiftung Research Grant via the ETH Zurich Foundation, SBB (ETH Mobility Initiative Partner).